OpenStudy (anonymous):

Find the distance between the points (13, 20) and (18, 8). 7 11 13

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

Well, I knew the formula I'm just having a hard time. ): I hate math!

hartnn (hartnn):

well. lets start by taking difference of x-co-ordinates, 13-18 =... ?

OpenStudy (anonymous):

-5.

hartnn (hartnn):

according to the formula, you need to square it , so (-5)^2 =... ?

OpenStudy (anonymous):

25.

hartnn (hartnn):

what about difference in y-co-ordinates ?

OpenStudy (anonymous):

Grr, ugh! ): I don't know. Please don't judge me!

hartnn (hartnn):

its ok.... difference of y-co-ordinates = 20-8 = 12 its square will be 144 so, just solve, \(d=\sqrt{144+25}\) and you'll get your distance :)

OpenStudy (anonymous):

Just do it on graph paper. Do the normal equation to get the hypotenuse. You can draw a triangle to figure out your other two legs.

OpenStudy (anonymous):

If your question is solved then close it.